OpenStudy (anonymous):

Best way to convert 10cm to nm? I never really understood the unit conversions that had to deal with 10^2 and 10^9 and stuff. Can someone please explain. I'm not talking about dimensional analysis

OpenStudy (anonymous):

there are 100 (or 10^2) cm in a meter. IN the same way, there are 10^9 nanometers in a meters. So if I were you, I would convert from centimeters to meters and then meters to nanometers.

OpenStudy (anonymous):

10^-9 you mean? Sorry I wrote it wrong in the question. But can you give me a quick refresher? It's embarrassing for me to ask but I need to know

OpenStudy (anonymous):

Dont feel bad, I have to set up proportions if i lose touch with this stuff. i think it is 10^9. I remember because nanometers are much smaller than meters. You need a lot of then to make up a meter.

OpenStudy (anonymous):

Ok you have you have 10 cm. to find the amount of meters set up a proportion, 10cm/xmeters=100cm/1meter.

OpenStudy (anonymous):

so you get .1 meters. then to go to nanometers, you know there are 10^9nm per meter. So 10^9 nm/1meter = x nm/.1 meter. solve for x. you should get 100 million

OpenStudy (anonymous):

i hope that helps

OpenStudy (anonymous):

I know this way is effective . But I mean by just using like how to get from 50*10^-2 to nm which is 10^-9

OpenStudy (anonymous):

I mean by using scientific notation

OpenStudy (marco26):

\[10cm*\frac{10^-2m}{1cm}*\frac{1nm}{10^-9m}\]